A community plans to build a facility to convert solar radiation to electrical power. The community requires 2.80 MW of power, and the system to be installed has an efficiency of 30.0% (that is, 30.0% of the solar energy incident on the surface is converted to useful energy that can power the community). Assuming sunlight has a constant intensity of 1 160 W/m2, what must be the effective area of a perfectly absorbing surface used in such an installation?
_________ m2
Answer #1

Given, Electrical power required, PE = 2.8*106 W

Intensity, I = 1160 W/m2

efficiency, n = 30 % = 0.3

Let the area of the surface be 'A' m2

then solar power will be, PS = 1160*A W

efficiency is the fraction of solar energy that is converted to electrical energy,i.e,

PE = n*PS

putting all the values

2.8*106 = 0.3*1160*A

A = 8045.98 m2